OpenStudy (anonymous):

Solve each system of equations by substitution -3v+3x=12 3v-6x=-21

OpenStudy (tkhunny):

1) Divide EVERYTHING by 3.

OpenStudy (skullpatrol):

$$-3v+3x=12$$ $$3v-6x=-21$$ $$\frac{1}{3}(-3v+3x)=\frac{1}{3}(12)$$ $$ \frac{1}{3} (3v-6x)=\frac{1}{3}(-21)$$

OpenStudy (skullpatrol):

Dividing by 3 is the same as multiplying by 1/3 right? @overtonka

OpenStudy (anonymous):

I really don't know how to do this o.o

OpenStudy (skullpatrol):

The multiplication property of equality allows you to multiply both sides of an equation by the same nonzero number and get an equivalent equation. Have you heard of this property in class?

OpenStudy (skullpatrol):

@overtonka

OpenStudy (anonymous):

No I havent

OpenStudy (skullpatrol):

Let me give you an example let's say 3=3, now I multiply both sides by 2 what would I get?

OpenStudy (skullpatrol):

@overtonka

OpenStudy (anonymous):

do u want the answer or how to get the answer

OpenStudy (ryan123345):

To atleast make the question easier divide everything by 3

OpenStudy (anonymous):

I tried that

OpenStudy (ryan123345):